# EMIT - Event-Based Masked Auto Encoding for Irregular Time Series

### v1

@@ I Introduction
To address these challenges, we can leverage a key characteristic of irregular time series: the rate of change as shown in Fig. 2. Understanding how data points change over time provides valuable insights. Therefore, this paper proposes a new pretraining framework with a masked autoencoder specifically designed for irregular time series based on the rate of change. Our approach masks points with a higher rate of change, which typically indicate significant or noteworthy events in the time series, and then attempts to reconstruct them. Instead of masking raw inputs, we mask the embeddings of these elements (time, value, feature) and aim to reconstruct the embeddings instead of the outputs. Masking the embeddings rather than raw data allows the model to develop features that are inherently robust and generalized. Our framework is versatile and can be integrated with any encoder that explicitly handles time, features, and values. Furthermore, we demonstrate through experiments on two benchmark datasets (Mimic-III [23] and PhysioNet Challenge 2012 [16] that the high-quality features learned through our pretraining task can be effectively fine-tuned with a limited set of labelled data. Our main contributions are summarized below:
@@
  • This paper proposes an algorithm to create masks based on the rate of change.
@@
Figure 2: An illustration of significant and insignificant events in the context of irregular time series. Points exhibiting a large rate of change are highlighted in green and are considered significant events. Conversely, points highlighted in orange have a relatively low rate of change and are considered insignificant events. Our model, EMIT, benefits from masking and reconstructing more points from the green section and fewer from the orange section.
@@ II-B1 Self-supervised learning in general machine learning
Supervised representation learning often requires manual labelling, which is time-consuming and expensive [11, 22]. To address these challenges, Self-Supervised Learning (SSL) became popular where a part of input was used to provide supervision signals. SSL observed a big success in CV [20, 19, 5, 17, 3] and NLP [36, 12, 28], which inspired its adoption to other domains like time series.
@@ II-B2 Self-supervised learning in regular time series
Regular time series, characterized by consistent intervals between measurements, benefit from various self-supervised learning (SSL) techniques, as detailed in surveys by [49, 30]. TimeNET [31] leverages a masked autoencoder to learn dense vector representations from time series of varying lengths. Triplet loss with time-based negative sampling is introduced by [15]. TST [48] employs a BERT-like [12] strategy using transformers to reconstruct masked values, replacing input values with zeros. TNC [44] applies a contrastive objective inspired by positive unlabelled learning, differentiating neighboring from non-neighboring samples. TS-TCC [14] utilizes dual contrastive mechanisms—temporal and contextual—to develop robust and discriminative representations. TARNet [9] alternates between pretext and downstream tasks, using task-informed data masking. Multimodal clinical time series contrastive learning with structured and physiological data is explored in [37]. BTSF [47] employs unique augmentation and bilinear fusion techniques to enhance feature representations by integrating time and spectral data. TF-C [50] focuses on integrating time-frequency components for SSL. SimMTM [13] proposes a framework inspired by manifold learning to reconstruct the original series through the weighted aggregation of neighboring masked series. FOCAL [27] extracts shared and distinct features from different modalities for robust SSL. InfoTS [29] uses a meta learner to select augmentations based on fidelity and variability. MHCCL [32] enhances contrastive learning with bidirectional masking in hierarchical clustering partitions. BasisFormer [33] adapts contrastive learning to extract basis from historical and future sections of a time series. Lastly, COMET [46] implements hierarchical contrasting at various levels within medical time series, from patient to observation.
@@ II-B3 Self-supervised learning in irregular time series
Researchers have explored several self-supervised strategies for irregular time series. These include window-based forecasting, contrastive learning, and masking-based reconstruction. STraTS [43] used forecasting to pretrain their model, leading to improved performance compared to models without a pretext task. PrimeNET [8] employed two time-sensitive pretext tasks: time-aware augmentations for contrastive learning and constant time masking to maintain the density of observations. In this paper, we leverage a key characteristic of irregular time series, the rate of change, to create a more effective masking strategy.
@@ III-A Problem Formulation
Problem: The primary objective is to learn a model $P$ from the given unlabelled pretraining dataset $D_{pt}$ containing $N$ sequences. The model should generate generalizable representations for specific downstream tasks. Our framework, EMIT, can be implemented with any architecture that explicitly encodes temporal, feature, and value information. In this work, we employ an architecture inspired by STraTS [43] as our backbone model.
@@ III-B2 Transformer Encoder
$\{\tilde{\mathbf{e}}_{1},\ldots,\tilde{\mathbf{e}}_{n}\}=\text{TransformerEncoder}(\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}),$ (7)
@@ III-B3 Time Series Embedding
where $\mathbf{W}_{a}\in\mathbb{R}^{d_{a}\times d}$, $\mathbf{b}_{a}\in\mathbb{R}^{d_{a}}$, and $\mathbf{u}_{a}\in\mathbb{R}^{d}$.
@@ III-C2 Applying Masks to Embeddings
$\{\tilde{\mathbf{e}}_{1},\ldots,\mathbf{\tilde{e}}_{\text{i, mask}},\tilde{\mathbf{e}}_{n}\}=\text{Encoder}(\{\mathbf{e}_{1},\ldots,\mathbf{e}_{\text{i, mask}},\mathbf{e}_{n}\})$ (18)
@@ IV-A1 Datasets
We assessed our proposed framework using two clinical datasets. Initially, we pretrained our model, followed by fine-tuning with various fractions of labelled data for the mortality prediction task. We adopted a similar methodology to the STraTS model for preprocessing and preparing the datasets to ensure a fair comparison. The MIMIC-III dataset, which is publicly available, includes electronic health records from over 40,000 patients in critical care units at Beth Israel Deaconess Medical Center, spanning from 2001 to 2012. The PhysioNet Challenge 2012111https://physionet.org/content/challenge-2012/1.0.0/ [16] dataset contains EHRs from more than 10,000 ICU stays and was used for a mortality prediction challenge in 2012. We excluded demographic information from our analysis to avoid potential biases and inaccuracies that could arise from imputation, given that these details might not be available for all patients. Table I highlights key statistics for both the datasets.
TABLE I: Dataset statistics.
@@ IV-A2 Metrics
TABLE II: Overall performance on MIMIC-III and PhysioNet-2012 (%). Bold numbers indicate the best performance of all methods.
Some baselines results are adopted from [43] due to similar experimental setup.
@@ IV-A4 Implementation
For both datasets, a Transformer Encoder [45] was used for encoding. The hyperparameters of the Transformer model were kept similar to STraTS for fair evaluation. We used 2 encoder blocks with 4 attention heads. Our scientific hyperparameters - event mask thresholds were chosen from {1e-1, 1e-2, 1e-3}, and the probability for masking insignificant events was chosen from {0, 0.1, 0.2, 0.3, 0.4, 0.5, 0.6, 0.7, 0.8, 0.9, 1.0}.
In EMIT pretraining, for the MIMIC dataset, the batch size is {256}, while for PhysioNet it is {128}, learning rate is {5e-4}, and dropout rates is {0.2} [42]. Using Adam [24] as the default optimizer, the learning rate weight decay value is set to {0}. For MIMIC, the time series length is 880 and the embedding dimension is 50. For PhysioNet, the maximum time series length is 791 and the embedding dimension is also 50. The maximum length is chosen as advised by STraTS. The models are pretrained for 100 epochs with early-stop. All experiments are conducted on a single GPU, from H100, A100, V100, and L40, with the training time per run varied from 4 to 7 hours.
For fine-tuning, for both datasets, the batch size is {32}, the learning rate is {5e-5}, the dropout rates are {0.2, 0.4}, and the weight decay values are {0, 1e-4}. The models are fine-tuned for 100 epochs with early-stop. The fine-tuning time per model varies from 1 to 3 hours. These experiments were repeated 5 times.
@@
Figure 5: Mortality prediction performance of different baseline models on PhysioNet 2012 dataset trained at different percentage of labelled data.
@@ IV-B Overall Performance (RQ1)
To evaluate the overall performance of our proposed method, EMIT, we conducted experiments on two healthcare datasets: MIMIC-III and PhysioNet-2012. We compared EMIT against several baseline models, including RNN-based models (GRU, GRU-D, InterpNet), convolution-based models (TCN), and transformer-based models (SaND, SeFT, STraTS). The evaluation metrics used were ROC-AUC, PR-AUC, and the minimum of recall and precision (min(Re, Pr)) across all possible decision thresholds.
TABLE III: Ablation studies on MIMIC-III and PhysioNet-201. Bold Numbers indicate the highest value for a given metric.
@@ IV-C Ablation Studies (RQ2 & RQ3)
Figure 6: Effect of different rate of change threshold on prediction performance using PhysioNet 2012 dataset.
Figure 7: Effect of different mask probabilities of insignificant events on prediction performance using PhysioNet 2012 dataset.

### v2

@@ I Introduction
To address these challenges, we can leverage a key characteristic of irregular time series: the rate of change as shown in Fig. 2. Understanding how data points change over time provides valuable insights. Therefore, this paper proposes a new pretraining framework with a masked autoencoder specifically designed for irregular time series based on the rate of change. Our approach masks points with a higher rate of change, which typically indicate significant or noteworthy events in the time series, and then attempts to reconstruct them. Instead of masking raw inputs, we mask the embeddings of these elements (time, value, feature) and aim to reconstruct the embeddings. Masking the embeddings rather than raw data allows the model to develop features that are inherently robust and generalized. Our framework is versatile and can be integrated with any encoder that explicitly handles time, features, and values. Furthermore, we demonstrate through experiments on two benchmark datasets (Mimic-III [23] and PhysioNet Challenge 2012 [16] that the high-quality features learned through our pretraining task can be effectively fine-tuned with a limited set of labelled data. Our main contributions are summarized below:
@@
  • This paper proposes an algorithm to create masks based on the rate of change in variables over time.
@@
Figure 2: An illustration of significant and insignificant events in the context of irregular time series. Points exhibiting a large rate of change are highlighted in green and are considered significant events. Conversely, points highlighted in orange have a relatively low rate of change and are considered insignificant events. Our model, EMIT, prioritizes masking and reconstruction of points associated with significant events, focusing on green regions that exhibit a large rate of change.
@@ II-B1 Self-supervised learning in general machine learning
Supervised representation learning often requires manual labelling, which is time-consuming and expensive [11, 22]. To address these challenges, Self-Supervised Learning (SSL) became popular where a part of input was used to provide supervisory signals. SSL observed a big success in CV [20, 19, 5, 17, 3] and NLP [36, 12, 28], which inspired its adoption to other domains like time series.
@@ II-B2 Self-supervised learning in regular time series
Regular time series, characterized by consistent intervals between measurements, benefit from various SSL techniques, as detailed in surveys by [49, 30]. TimeNET [31] leverages a masked autoencoder to learn dense vector representations from time series of varying lengths. Triplet loss with time-based negative sampling is introduced by [15]. TST [48] employs a BERT-like [12] strategy using transformers to reconstruct masked values, replacing input values with zeros. TNC [44] applies a contrastive objective inspired by positive unlabelled learning, differentiating neighboring from non-neighboring samples. TS-TCC [14] utilizes dual contrastive mechanisms—temporal and contextual—to develop robust and discriminative representations. TARNet [9] alternates between pretext and downstream tasks, using task-informed data masking. Multimodal clinical time series contrastive learning with structured and physiological data is explored in [37]. BTSF [47] employs unique augmentation and bilinear fusion techniques to enhance feature representations by integrating time and spectral data. TF-C [50] focuses on integrating time-frequency components for SSL. SimMTM [13] proposes a framework inspired by manifold learning to reconstruct the original series through the weighted aggregation of neighboring masked series. FOCAL [27] extracts shared and distinct features from different modalities for robust SSL. InfoTS [29] uses a meta learner to select augmentations based on fidelity and variability. MHCCL [32] enhances contrastive learning with bidirectional masking in hierarchical clustering partitions. BasisFormer [33] adapts contrastive learning to extract basis from historical and future sections of a time series. Lastly, COMET [46] implements hierarchical contrasting at various levels within medical time series, from patient to observation.
@@ II-B3 Self-supervised learning in irregular time series
Researchers have explored several self-supervised strategies for irregular time series. These include window-based forecasting, contrastive learning, and masking-based reconstruction. STraTS [43] used forecasting to pretrain their model, leading to improved performance compared to models without a pretext task. PrimeNET [8] employed two time-sensitive pretext tasks: time-aware augmentations for contrastive learning and constant time masking. In this paper, we leverage a key characteristic of irregular time series, the rate of change, to create a more effective masking strategy.
@@ III-A Problem Formulation
Problem: The primary objective is to learn a model $P$ from the given unlabelled pretraining dataset $D_{pt}$ containing $N$ sequences. The model should generate generalizable representations for specific downstream tasks.
Our framework, EMIT, can be implemented with any architecture that explicitly encodes temporal, feature, and value information. In this work, we employ an architecture inspired by STraTS [43] as our backbone model.
@@ III-B2 Transformer Encoder
$\{\tilde{\mathbf{e}}_{1},\ldots,\tilde{\mathbf{e}}_{n}\}=\text{TrEnc}(\{\mathbf{e}_{1},\ldots,\mathbf{e}_{n}\}),$ (7)
@@ III-B3 Time Series Embedding
where $\mathbf{W}_{a}\in\mathbb{R}^{d_{a}\times d}$, $\mathbf{b}_{a}\in\mathbb{R}^{d_{a}}$, and $\mathbf{u}_{a}\in\mathbb{R}^{d_{a}}$.
@@ III-C2 Applying Masks to Embeddings
$\{\tilde{\mathbf{e}}_{1},\ldots,\tilde{\mathbf{e}}_{\text{i, mask}},\ldots,\tilde{\mathbf{e}}_{n}\}$ $=$ (18)
$\operatorname{TrEnc}(\{\mathbf{e}_{1},\ldots,\mathbf{e}_{\text{i,mask}},\ldots,\mathbf{e}_{n}\})$
@@ IV-A1 Datasets
We assessed our proposed framework using two clinical datasets. Initially, we pretrained our model, followed by fine-tuning with various fractions of labelled data for the mortality prediction task. We adopted a similar methodology to the STraTS[43] for preprocessing and preparing the datasets to ensure a fair comparison. The MIMIC-III dataset, which is publicly available, includes electronic health records from over 40,000 patients in critical care units at Beth Israel Deaconess Medical Center, spanning from 2001 to 2012. The PhysioNet Challenge 2012111https://physionet.org/content/challenge-2012/1.0.0/ [16] dataset contains EHRs from more than 10,000 ICU stays and was used for a mortality prediction challenge in 2012. We excluded demographic information from our analysis to avoid potential biases and inaccuracies that could arise from imputation, given that these details might not be available for all patients. Table I highlights key statistics for both the datasets.
TABLE I: Dataset statistics. Adapted from [43]
@@ IV-A2 Metrics
TABLE II: Overall performance on MIMIC-III and PhysioNet-2012 on 50% labelled data. Bold numbers indicate the best performance of all methods. Some baselines results are adopted from [43] due to similar experimental setup.
@@ IV-A4 Implementation
For both datasets, a Transformer Encoder [45] was used for encoding. For fair evaluation, the hyperparameters of the Transformer model were kept similar to STraTS[43]: 2 encoder blocks and 4 attention heads. Our scientific hyperparameters - rate of change threshold was chosen from {1e-1, 1e-2, 1e-3}, and the probability for masking insignificant events was chosen from {0, 0.1, …, 1.0}.
In EMIT pretraining, the MIMIC dataset used a batch size of 256 and the PhysioNet dataset used 128. Both datasets had a learning rate of 5e-4 and a dropout rate of 0.2[42]. The Adam optimizer[24] was used with a weight decay of 0. For MIMIC, the time series length was 880 with an embedding dimension of 50; for PhysioNet, the maximum time series length was 791, also with an embedding dimension of 50, as advised by STraTS[43]. Models were pretrained for 100 epochs with early stopping. All experiments were conducted on a single GPU from H100, A100, V100, or L40, with training times ranging from 4 to 7 hours.
For fine-tuning, both datasets used a batch size of 32, a learning rate of 5e-5, dropout rates of {0.2, 0.4}, and weight decay values of {0, 1e-4}. Models were fine-tuned for 100 epochs with early stopping, and the fine-tuning time per model ranged from 1 to 3 hours. These experiments were repeated 5 times.
@@
Figure 5: Mortality prediction performance of different baseline models on PhysioNet Challenge 2012 dataset trained at different percentage of labelled data.
@@ IV-B Overall Performance (RQ1)
To evaluate the overall performance of our proposed method, EMIT, we conducted experiments on two healthcare datasets: MIMIC-III and PhysioNet-2012. We compared EMIT against several baseline models, including RNN-based models [10, 4, 40], convolution-based models[1] and transformer-based models [41, 21, 43]. The evaluation metrics used were ROC-AUC, PR-AUC, and the minimum of recall and precision (min(Re, Pr)) across all possible decision thresholds.
TABLE III: Ablation studies on MIMIC-III and PhysioNet-2012. Bold Numbers indicate the highest value for a given metric.
@@ IV-C Ablation Studies (RQ2 & RQ3)
Figure 6: Effect of different rate of change threshold on prediction performance using PhysioNet Challenge 2012 dataset.
Figure 7: Effect of different mask probabilities of insignificant events on prediction performance using PhysioNet Challenge 2012 dataset.